The rate constant for this second‑order reaction is 0.560 M−1⋅s−10.560 M−1⋅s−1 at 300 ∘C.300 ∘C.

A⟶productsA⟶products

How long, in seconds, would it take for the concentration of AA to decrease from 0.910 M0.910 M to 0.320 M?0.320 M?
Answer #1

- a- 1 FC Co kt C = final concentration Cocinitial concentration. I - I = 0.560xt 0.320 0.910 3. 126 - 1.098 & 0.560xt 2.02 =